Calculate the average annual growth rates of the following countries below:

Country 1980 2014
United States 29288 51958
Japan 19147 35574
Mexico 11954 15521
Indonesia 2249 9797

How long will it take the countries above to double their GDPs?
Answer #1

We use the rule that future value = present value * (1 + rate of growth)^time

For example, average annual growth rate of Mexico is (15521/11954)^(1/(2014 - 1980)) - 1 = 0.77% per year

Doubling time = rule of 72

= 72/ growth rate

For Mexico, the doubling time for GDP = 72/0.77 = 94 years approximately.

The same for other countries is done below.

Country 1980 2014 Rule (FV = PV(1 + r)^n) Average annual growth Doubling Time (years)
United States 29288 51958 (GDP 2014/GDP 1980)^(1/(2014-1980))-1 1.70 42.34
Japan 19147 35574 (GDP 2014/GDP 1980)^(2014-1980)-1 1.84 39.16
Mexico 11954 15521 (GDP 2014/GDP 1980)^(2014-1980)-1 0.77 93.39
Indonesia 2249 9797 (GDP 2014/GDP 1980)^(2014-1980)-1 4.42 16.28